# On the optimal prediction of extreme events in heavy-tailed time series with applications to solar flare forecasting

### v1

@@ 3.2.2 Asymptotic Optimality of the Estimated Predictor
$|\widehat{U}_{t+h}-U_{t+h}^{*}|\leq|\widehat{U}_{t+h}-\widetilde{U}_{t+h}|+|\widetilde{U}_{t+h}-U_{t+h}^{*}|.$ (3.18)
$f_{j}(x_{0},x_{1},\cdots):=\sum_{k=0}^{\infty}a_{k}(\phi)x_{j+k},$
where $Y_{t}=\sum_{k=0}^{\infty}a_{k}(\phi)\epsilon_{t-k}$ is the causal moving average representation of the one-dimensional AR$(d)$ time-series. Now, the ergodicity of the map $T$ entails the ergodicity of the process $\{Z_{t}=Z_{0}\circ T^{t}\}$. In order to use Theorem 3.1 from Adams & Nobel (2012), we must also have that the collection of all half-spaces $H_{\varphi,\,\tau}$ is a separable Vapnik-Chervonenkis class, this is shown in Lemma A.6 in the Appendix. The latter uniform law of large numbers theorem now yields
which shows that the first term on the right side in (3.18) vanishes.
Now we handle the second term on the right side in (3.18). Recall that $Z$ in (3.15) is chosen to be independent from $(Y_{t:(t-d+1)},\widehat{\phi})$. Thus, $Z$ is independent from $\widehat{Y}_{t+h}(\widehat{\phi})=\widehat{\phi}(h)^{\top}Y_{t:(t-d+1)}$ and $\widehat{\phi}(h)$ and in view of (3.15), we can write:
$\widetilde{U}_{t+h}-U_{t+h}^{*}=\mathbb{E}\left[\mathbb{I}(\widehat{\phi}(h)^{\top}Z\leq\widehat{Y}_{t+h}(\widehat{\phi}))-\mathbb{I}(\phi(h)^{\top}Z\leq\widehat{Y}_{t+h}(\phi))\Big{|}\widehat{\phi},Y_{t:(t-d+1)}\right]\$ (3.19)
Introduce
In view of (3.19),
$\leq\mathbb{E}|\mathbb{I}[\xi_{n}\leq 0]-\mathbb{I}[\xi\leq 0]|$
[…]
This completes the proof. ∎
@@ 3.3 The Optimal Extremal Precision for MA($\infty$) Models
Let $a,b\in\mathcal{S}_{\alpha}$. Assume that $p_{\xi(a)},p_{\xi(b)}>0$. Then
@@
The process defined by (3.22) is invertible, i.e., there exists a real sequence $\{b_{j}\}_{j=0}^{\infty}$ such that for all $t$,
$\epsilon_{t}=\sum_{j=0}^{\infty}b_{j}Y_{t-j}$
almost surely.
@@
$\lambda_{h}^{\rm(opt)}$ $=\sum_{j=0}^{\infty}\kappa_{+,+}(a_{j+h},a_{j+h})\left(\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{h})}\right)$
$=\sum_{j=0}^{\infty}\kappa_{+}(a_{j+h})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}=\frac{\eta_{+}({a},{h})}{\eta_{+}({a},{0})}.$ (3.30)
At time $t$, the data available for prediction of $Y_{t+h}$ consists of at least $Y_{t}$. One may wonder how $Y_{t}$ compares as a predictor to the optimal predictor. Proposition 3.2 describes circumstances under which these two predictors have the same asymptotic precision, or the optimal asymptotic precision can at least be expressed or bounded in terms of the asymptotic precision of ±$\pm Y_{t}$.
@@
(iii) If $p_{\epsilon}=1/2$, we have $\eta_{-}({a},{0})=\eta_{+}({a},{0})$ and
$\lambda(Y_{t+h},Y_{t})+\lambda(Y_{t+h},-Y_{t})=\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}}).$
Proof of Proposition 3.2.
Recall from (3.30) that
$\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}})=\sum_{j\geq 0}\kappa_{+}(a_{j+h})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}.$ (3.34)
On the other hand, by Lemma 3.4,
$\lambda(Y_{t+h},Y_{t})$ $=\sum_{j\geq 0}\kappa_{+,+}(a_{j+h},a_{j})\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\Big{)}$ (3.35)
$=\sum_{j\geq 0}\kappa_{+,+}(a_{j+h},a_{j})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})},$ (3.36)
[…]
$=\kappa_{+}(a_{j+h}),$
and hence the right-hand sides of (3.34) and (3.36) are identical. This proves (3.32).
On the other hand, if
$\lambda(Y_{t+h},-Y_{t})$ $=\sum_{j\geq 0}\kappa_{+,-}(a_{j+h},a_{j})\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{-}({a},{0})}\Big{)},$ (3.37)
For all $j\geq 0$, we have
$\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{+}({a},{0})}\Big{)}\leq\Big{(}1\bigvee\frac{\eta_{-}({a},{0})}{\eta_{+}({a},{0})}\Big{)}\cdot\Big{(}\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j}|^{\alpha}}{\eta_{-}({a},{0})}\Big{)},$ (3.38)
where we used the fact that $x\wedge(wy)\leq(1\vee w)(x\wedge y)$, for all non-negative $x,y,$ and $w$. At the same time, for the sum of the weights in the right-hand sides of (3.35) and (3.37), we have
$=\kappa_{+}(a_{j+h}),$ (3.39)
[…]
By (3.31), however, we have $|a_{j+h}|^{\alpha}\wedge|a_{j}|^{\alpha}=|a_{j+h}|^{\alpha}$, and in view of (3.34) the left-hand side of the last relation equals $\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}})$. This completes the proof of the upper bound in (3.33). In view of (3.34), and Relations (3.35), (3.37), and (3.3), the lower bound in (3.33) is an immediate consequence of the inequalities:
@@
4 Simulations
We illustrate several results and ideas discussed in previous sections with simulation studies.
4.1 Autoregressive Models
Optimal extreme event prediction for autoregressive models was discussed in Section 3.1. We restate the model from that section for the sake of convenience:
$Y_{t}=\sum_{i=1}^{d}\phi_{i}Y_{t-i}+\epsilon_{t},\ \ \ t\in{\mathbb{Z}}.$ (4.1)
It was shown that for a stationary process satisfying (4.1), for $h\geq 1$, the optimal predictor of $\mathbb{I}(Y_{t+h}>y_{0})$ is $\mathbb{I}(\widehat{Y}_{t+h}(\phi)>\tau)$ for an appropriate choice of $\tau$ (Theorem 3.2). In Section 3.2.1, we discussed prediction of $\mathbb{I}(Y_{t+h}>F_{Y}^{\leftarrow}(p))$, $F_{Y}^{\leftarrow}(p)$ being the $p$th quantile of the marginal distribution of $Y_{t}$; we considered the use of the predictor $\mathbb{I}(\widehat{Y}_{t+h}(\widehat{\phi})\geq\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$ as a practical alternative to the predictor $\mathbb{I}(\widehat{Y}_{t+h}(\widehat{\phi})\geq F_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$, which we shall call the oracle predictor.
The most straightforward way to compute the estimate $\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ of $F_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ is to use an empirical quantile, as was done in Section 3.2.1. When $\{Y_{t}\}$ is a heavy-tailed process and $p$ is close to one, an alternative is to use an extreme quantile estimator based on extreme value theory (see, e.g., Weissman, 1978; Drees, 2003; Chavez-Demoulin & Guillou, 2018, and the references therein).
Figure 4.1 shows the results of a simulation study that compared the performance of the oracle predictor, a predictor based on the empirical quantile estimator, and a predictor based on an extreme quantile estimator. Full details on the design of the study can be found in Section A.3.1. Data was generated from an AR(5) model with $t_{1}$ innovations. A large number of training set-test set pairs were generated; an AR(5) model was fit to each training set, and then predictions were made on the test set for each $p\in\{0.90,0.95,0.99,0.999\}$ using each predictor.
Orange boxplots represent the precisions of the oracle predictor. As expected, as $p$ approaches one, the oracle precision converges to the asymptotic oracle precision computed using Theorem 3.4, which is represented by the dashed orange line. The blue boxplots represent the precisions of the predictor based on the empirical quantile estimator; this predictor exhibits strong performance, as its median precisions are the same as the corresponding oracle precisions. However, this is not the case when $p=0.999$; the median precision is noticeably smaller than the oracle precision. This is not surprising, as the training set size was $10,000$. The black boxplots represent precisions of the predictor based on the extreme quantile estimator. Interestingly, these lie well above the solid line, so that predictor performed better than the oracle predictor. It turned out that the extreme quantile estimates were substantially too large, so that predictor was calibrated at a higher level than the other predictors. This is not guaranteed to yield higher precisions, but it ended up doing so in this case.
Figure 4.1: Results from the autoregressive model simulation study. Data was generated from a particular AR(5) model. One hundred training set-test set pairs were generated; the model was fit to the training set, and the fitted model was used to predict exceedances of high quantiles in the test set. The quantile levels used are on the horizontal axis. On each of the 100 runs, the precision was calculated from the test set predictions. The dashed orange line represents the asymptotic precision of the optimal predictor. See Section A.3.1 for full information about the study design.
4.2 Infinite-Variance FARIMA Models
We next illustrate optimal extreme event prediction with an interesting model that is both long-range dependent and heavy-tailed, namely, a fractional autoregressive integrated moving average (FARIMA) model. The idea to consider FARIMA came from our reading of Stanislavsky et al. (2009), where such models were applied to the GOES X-ray flux data. In Section 5, we fit FARIMA models to the same data and use the fitted models to predict the high fluxes that define solar flares. In this section, we merely illustrate Theorem 3.4 in the context of FARIMA models for the special case of symmetric $\alpha$-stable innovations.
A FARIMA($p$, $d$, $q$) model is defined by the equation
@@
for some $c>0$ (Samorodnitsky & Taqqu, 1994; Nolan, 2020). Therefore, the innovations as well as the resulting FARIMA time series will have infinite variances and in fact infinite means for $0<\alpha\leq 1$. One can consider FARIMA models with asymmetric regularly varying innovations in the domain of attraction of $\alpha$-stable laws (see, e.g., Section 5 below and Kokoszka & Taqqu, 1996).
For the sake of simplicity, we consider FARIMA$(0,d,0)$ models, where (4.2) becomes
@@
It follows from Theorem 2.1 in Kokoszka & Taqqu (1995) that if $d$ is non-integral, $\alpha\in(1,2)$, and $d<1-1/\alpha$, then the unique solution to (4.4) is
$Y_{t}=\sum_{j=0}^{\infty}a_{j}\epsilon_{t-j},$ (4.5)
where the terms of $a=\{a_{j}\}$ are given by
@@
$\Gamma$ being the gamma function.
It is well-known that for $d>0,$
@@
(see, e.g., Lemma 3.1 in Kokoszka & Taqqu, 1995). This implies that $a=\{a_{j}\}$ belongs to $\mathcal{S}_{\alpha}$ provided $d<1-1/\alpha$. By (4.3) and Lemma 3.3 it follows that the solution in (4.5) is finite almost surely, whereas Theorem 2.3 of Kokoszka & Taqqu (1995) implies that the solution is invertible.
By Theorem 3.1, for any $p\in(0,1)$ and $y_{0}\in\mathbb{R}$ such that $\mathbb{P}(Y_{t+h}\leq y_{0})=p$, the optimal predictor of $\mathbb{I}(Y_{t+h}>y_{0})$ is the indicator of whether $\sum_{j=0}^{\infty}a_{j+h}\epsilon_{t-j}$ is greater than $\tau$ for a suitably chosen constant $\tau$. Theorem 3.4 tells us that the asymptotic precision of the optimal predictor is the ratio of series in (3.29). For various pairs of $d$ and $\alpha$ satisfying the constraints $\alpha\in(1,2)$, $d\in(0,1-1/\alpha)$, we approximated the asymptotic precision of the optimal predictor by truncating the series in (3.29). The approximations for one-step-ahead prediction are shown in Figure 4.2 (left).
We see that for a fixed $\alpha$, as $d$ increases, the optimal asymptotic precision increases. This is intuitive as the level of long-range dependence in $\{Y_{t}\}$ increases as $d$ increases. For example, when $d=0$, (4.4) simplifies to $Y_{t}=\epsilon_{t}$, $t\in{\mathbb{Z}}$, so $\{Y_{t}\}$ cannot be long-range dependent as it consists of independent random variables. The higher the level of long-range dependence, the more $Y_{t+h}$ depends on $Y_{s}$’s in the distant past, and the better the optimal predictor performs. On the other hand, for a fixed $d$, the optimal asymptotic precision decreases as $\alpha$ increases.
Figure 4.2: Left panel: a contour plot showing how the one-step-ahead asymptotic precision of the optimal predictor for the FARIMA(0, $d$, 0) model (4.4) varies as $\alpha$ and $d$ vary. Right panel: the asymptotic optimal precision as a function of $h$ for the combination of $\alpha=1.4$ and $d=0.19$ indicated by the red dot in the left panel.
5 Data Analysis
Figure 5.1: A plot of the combined time series from the maxima of solar cycles 23 and 24 corresponding to date ranges 1/1/2000-12/31/2002 and 9/1/2011-5/31/2014, respectively. The horizontal lines correspond to the 0.9, 0.95, and 0.99 sample quantiles of the combined dataset. Note that the time series is plotted on a logarithmic scale.
@@ 5 Data Analysis
We processed the downloaded data as follows. First, we only retained measurements recorded between January 2000 and December 2002 or between September 2011 and May 2014. These two periods roughly represent the maxima of solar cycles 23 and 24, respectively; the flux and flaring activity are elevated during the maximum of a solar cycle. These periods were identified by looking for the peak periods in the sunspot time series displayed in the U.S. Space Weather Prediction Center’s Solar Cycle Progression product [^1] Available at: https://www.swpc.noaa.gov/products/solar-cycle-progression. To smooth the data, we took the maximum flux over each hour of the day. This produced a time series of length 50,400. Next, we used linear interpolation to impute a small number of missing values, which constituted a little less than 1% of the values. The final time series is displayed in Figure 5.1. Hourly maxima are colored based on which solar cycle they belong to. Note that since we have concatenated two time series separated by several years, observation indices instead of times are shown on the horizontal axis. Figure 5.1 suggests that it is reasonable to treat the concatenated series as a single time series.
We divided the final time series into windows of size 4,320, which is the number of hours in 180 days. Consecutive windows differed in their starting times by twelve observations, so there were 3,840 windows. Over each window, we fit three different kinds of models, and predicted exceedances of several thresholds at several lead times.
The first kind of model, which we shall call the baseline model, predicts that a future value $X_{t+h}$ will exceed a threshold if and only if the most recent observation $X_{t}$ exceeds it. The second kind of model is an AR($d$) model, which we fit using least squares. Predictions were made using the methodology described in Algorithm 1 in Section 3.2.1.
The third kind of model is a FARIMA$(0,d,0)$ model with symmetric $\alpha$-stable innovations. See Section 4.2 for the definition of a FARIMA model. This choice of model was inspired by Stanislavsky et al. (2009), which found that the flux time series can be modeled reasonably well using a FARIMA$(0,d,0)$ model. As in Section 4.2, $d$ is non-integral, $\alpha\in(1,2)$, and $d<1-1/\alpha$. Under the model, the centered observations $Y_{t}$ satisfy
$Y_{t}=\sum_{j=0}^{\infty}a_{j}\epsilon_{t-j},$ (5.1)
where the $a_{j}$’s are defined by the recurrence relation in (4.6). Also, $\{Y_{t}\}$ is stationary, so the $Y_{t}$’s are identically distributed. Note that by Lemma 3.3, the $Y_{t}$’s are regularly varying with tail index $\alpha$.
Since the observations are maxima over blocks of a relatively large number of observations, sixty, we can approximate the marginal distribution of the centered observations by a GEV (generalized extreme value) distribution. Because of the celebrated Fisher-Tippett-Gnedenko theorem and its extensions, a GEV distribution is a common choice of model for block maxima (Leadbetter et al., 1983; Embrechts et al., 1997; Beirlant et al., 2004). The GEV distribution with location parameter $\mu$, scale parameter $\sigma>0$, and shape parameter $\xi$ has distribution function
@@
Figure 5.2: Left panel: estimates of $\alpha$ for the FARIMA$(0,d,0)$ models with symmetric $\alpha$-stable innovations that were fit over the training windows. Center panel: estimates of $d$ for the various models. Right panel: a scatterplot showing the relationship between $\widehat{\alpha}$ and $\widehat{d}$; for these models, $d$ must be less than $1-1/\alpha$.
The right panel of Figure 5.2 reveals some interesting aspects of the relationship between $\widehat{\alpha}$ and $\widehat{d}$. Recall that $d<1-1/\alpha$, where for many training windows, $\widehat{d}$ is as large as it is permitted to be, given $\widehat{\alpha}$. This suggests that over many windows, the time series exhibits very strong long-range dependence. However, this could also be partly due to presence of complex non-stationary trends in the data that can often be confounded with long-range dependence (Stoev et al., 2005).
Suppose that we have fit a FARIMA$(0,d,0)$ model to centered training observations $Y_{t},\ldots,Y_{t-n+1}$, with $\widehat{d}$ being the estimate of $d$. Also suppose that we are interested in forecasting exceedance of a threshold at time $t+h$ using the most recent $\ell$ observations, $Y_{t},\ldots,Y_{t-\ell+1}$. The training set size $n$ should be considerably larger than $\ell$. By Remark 3.2, we can do this by thresholding appropriately the quantity $\sum_{r=0}^{\ell-1}c_{r}Y_{t-r}$, where
@@
Table 5.1 summarizes the results from our prediction experiments. We sought to predict exceedances of the $p$th marginal quantile $h$ steps ahead, for $p=0.90,0.95,0.99$ and $h=1$ and $6$ (see also Table A.2 for $h=12$ and $18$). The number of observations used to computed predictions from FARIMA models was 168, the number of hours in one week; this value was also used as the order of the AR models that we fit. The True Skill Statistic (TSS) is commonly used to evaluate solar flare forecasting methods; see Remark A.6 for more information about it. Since the patterns in the TSS section of Table 5.1 are very similar to those in the precision section, we shall only comment on the latter. In each row, in the section for each metric, the largest number is in bold.

### v2

@@ 3.2.2 Asymptotic Optimality of the Estimated Predictor
See Section A.2.4. ∎
@@
Remark 3.6.
See Section A.2.6 for the results of a simulation study of the performance of the optimal AR predictor $\mathbb{I}(\widehat{Y}_{t+h}(\phi)>F_{\widehat{Y}_{t+h}(\phi)}^{\leftarrow}(p))$ and two approximations to it of the form $\mathbb{I}(\widehat{Y}_{t+h}(\widehat{\phi})\geq\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p))$, where the quantile estimator $\widehat{F}_{\widehat{Y}_{t+h}(\widehat{\phi})}^{\leftarrow}(p)$ is either a sample quantile as in this section or an estimator based on extreme value theory.
@@ 3.3 The Optimal Extremal Precision for MA($\infty$) Models
Let $a,b\in\mathcal{S}_{\alpha}$. Assume that $p_{\xi(a)}>0$ and $p_{\xi(b)}>0$. Then
@@
The process defined by (3.20) is invertible, i.e., there exists a real sequence $\{b_{j}\}_{j=0}^{\infty}$ such that for all $t$, $\epsilon_{t}=\sum_{j=0}^{\infty}b_{j}Y_{t-j}$ almost surely.
@@
$\lambda_{h}^{\rm(opt)}=\sum_{j=0}^{\infty}\kappa_{+,+}(a_{j+h},a_{j+h})\left(\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}\bigwedge\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{h})}\right)=\sum_{j=0}^{\infty}\kappa_{+}(a_{j+h})\frac{|a_{j+h}|^{\alpha}}{\eta_{+}({a},{0})}=\frac{\eta_{+}({a},{h})}{\eta_{+}({a},{0})}.$ (3.28)
At time $t$, the data available for prediction of $Y_{t+h}$ consists of at least $Y_{t}$. One may wonder how a predictor based solely on $Y_{t}$ compares to the optimal predictor. Proposition 3.2 describes circumstances under which these two predictors have the same extremal precision, or the extremal optimal precision can at least be expressed or bounded in terms of the extremal precision of ±$\pm Y_{t}$.
@@
(iii) If $p_{\epsilon}=1/2$, we have $\eta_{-}({a},{0})=\eta_{+}({a},{0})$ and $\lambda(Y_{t+h},Y_{t})+\lambda(Y_{t+h},-Y_{t})=\lambda(Y_{t+h},\widehat{Y}_{t+h}^{\text{(opt)}})$.
Proof.
See Section A.2.10. ∎
@@
4 Data Analysis
We illustrate the use of our methodological and theoretical results with an analysis of soft X-ray flux data. As stated in Section 1, the peak value of the flux during a flare determines the flare’s strength. We fit several models to a time series of flux values and used them to predict exceedances of high thresholds at multiple lead times. Section 4.1 describes how the raw data was processed into the data used to fit models. FARIMA models were one of the types of models we fit; Section 4.2 gives background on them, while Section 4.3 describes how they were fit and discusses the fitted models. The results for the various prediction scenarios are presented in Section 4.4.
4.1 Data Processing
[…]
We processed the downloaded data as follows. First, we only retained measurements recorded between January 2000 and December 2002 or between September 2011 and May 2014. These two periods roughly represent the maxima of solar cycles 23 and 24, respectively; the flux and flaring activity are elevated during the maximum of a solar cycle. These periods were identified by looking for the peak periods in the sunspot time series displayed in the U.S. Space Weather Prediction Center’s Solar Cycle Progression product. To smooth the data, we took the maximum flux over each hour of the day. This produced a time series of length 50,400. Next, we used linear interpolation to impute a small number of missing values, which constituted a little less than 1% of the values. The final time series is displayed in Figure 4.1. Hourly maxima are colored based on which solar cycle they belong to. Note that since we have concatenated two time series separated by several years, observation indices instead of times are shown on the horizontal axis. Figure 4.1 suggests that it is reasonable to treat the concatenated series as a single time series.
Figure 4.1: A plot of the combined soft X-ray flux time series from the maxima of solar cycles 23 and 24, which correspond to the date ranges 1 January 2000 to 31 December 2002 and 1 September 2011 to 31 May 2014, respectively. The horizontal lines correspond to the 0.9, 0.95, and 0.99 sample quantiles of the combined dataset. Note that the time series is plotted on a logarithmic scale.
We divided the final time series into windows of size 4,320, which is the number of hours in 180 days. Consecutive windows differed in their starting times by twelve observations, so there were 3,840 windows. Over each window, we fit three different kinds of models, and we predicted exceedances of several thresholds $h$ hours after the last hour in the window for several values of $h$. The first kind of model, which we shall call the baseline model, predicts that a future maximum flux will exceed a threshold if and only if the most recent maximum flux exceeds it. The second kind of model is an AR($d$) model, which we fit using least squares after centering the flux maxima in the training window; the maxima were centered because the flux is positive. We initially tried fitting AR models using LAD estimation, but found that the fitted models were not calibrated as well as models fit using least squares; see Table A.3 and Figure A.3 for more details. Predictions were made using the methodology described in Algorithm 1 in Section 3.2.1. The third kind of model is a FARIMA$(0,d,0)$ model with symmetric $\alpha$-stable innovations, which was also fit to centered data. We next give necessary background on FARIMA models.
4.2 Background on FARIMA Models
The idea to consider fractional autoregressive integrated moving average (FARIMA) models came from our reading of Stanislavsky et al. (2009), where such models were applied to the flux data. A FARIMA($p$, $d$, $q$) model is defined by the equation
@@ 4.2 Background on FARIMA Models
for some $c>0$ (Samorodnitsky & Taqqu, 1994; Nolan, 2020). Therefore, the innovations as well as the resulting FARIMA time series will have infinite variances and in fact infinite means for $0<\alpha\leq 1$. One can instead consider FARIMA models with asymmetric regularly varying innovations in the domain of attraction of $\alpha$-stable laws (see, e.g., Kokoszka & Taqqu, 1996).
For the sake of simplicity, we consider FARIMA$(0,d,0)$ models, for which (4.1) becomes
@@
Stanislavsky et al. (2009) found that the flux time series can be modeled reasonably well using such a model. It follows from Theorem 2.1 in Kokoszka & Taqqu (1995) that if $d$ is non-integral, $\alpha\in(1,2)$, and $d<1-1/\alpha$, then the unique solution to (4.3) is $Y_{t}=\sum_{j=0}^{\infty}a_{j}\epsilon_{t-j}$, where the terms of $a=\{a_{j}\}$ are given by
@@
$\Gamma$ being the gamma function. It is well-known that for $d>0$,
@@
(see, e.g., Lemma 3.1 in Kokoszka & Taqqu, 1995). This implies that $a=\{a_{j}\}$ belongs to $\mathcal{S}_{\alpha}$ provided that $d<1-1/\alpha$. It follows from (4.2) and Lemma 3.3 that the solution $\{Y_{t}\}$ to (4.3) is finite almost surely; Theorem 2.3 of Kokoszka & Taqqu (1995) implies that the solution is invertible.
Remark 4.1.
Under the assumptions about $d$ and $\alpha$ given above, for any $h\geq 1$, the extremal optimal precision $\lambda_{h}^{\rm(opt)}$ can be computed as the ratio $\eta_{+}({a},{h})/\eta_{+}({a},{0})$, according to Theorem 3.4. Given $d$, $\alpha$, and the extremal skewness $p_{\epsilon}$ of the innovations, $\lambda_{h}^{\rm(opt)}$ can be approximated by truncating the series defining $\eta_{+}({a},{h})$ and $\eta_{+}({a},{0})$ using some large number of initial terms. We used this approach to investigate how the extremal optimal precision for a FARIMA$(0,d,0)$ model varies with $d$ and $\alpha$; see Section A.3.1 for the results.
4.3 FARIMA Modeling of the Soft X-Ray Flux
@@
$Y_{t}=(1-B)^{-d}\epsilon_{t},\quad t\in{\mathbb{Z}},$
[…]
Since the uncentered observations are maxima over blocks of a relatively large number of values, sixty, we can approximate the marginal distribution of the centered observations by a GEV (generalized extreme value) distribution. Because of the celebrated Fisher-Tippett-Gnedenko theorem and its extensions, a GEV distribution is a common choice of model for block maxima (Leadbetter et al., 1983; Embrechts et al., 1997; Beirlant et al., 2004). The GEV distribution with location parameter $\mu$, scale parameter $\sigma>0$, and shape parameter $\xi$ has distribution function
@@
Figure 4.2: Left: estimates of $\alpha$ for the FARIMA$(0,d,0)$ models with symmetric $\alpha$-stable innovations that were fit over the training windows. Center: estimates of $d$ for the various models. Right: a scatterplot showing the relationship between $\widehat{\alpha}$ and $\widehat{d}$; for these models, $d$ must be less than $1-1/\alpha$, hence the curve in the upper left.
The right panel of Figure 4.2 reveals some interesting aspects of the relationship between $\widehat{\alpha}$ and $\widehat{d}$. Recall that $d<1-1/\alpha$; for many training windows, $\widehat{d}$ is as large as it is permitted to be, given $\widehat{\alpha}$. This suggests that over many windows, the time series exhibits very strong long-range dependence. However, this could also be partly due to the presence of complex non-stationary trends in the data that can often be confounded with long-range dependence (Stoev et al., 2005).
Suppose that we have fit a FARIMA$(0,d,0)$ model to centered training values $Y_{t},\ldots,Y_{t-n+1}$, with $\widehat{d}$ being the estimate of $d$. Also suppose that we are interested in forecasting exceedance of a threshold at time $t+h$ using the most recent $\ell$ observations, $Y_{t},\ldots,Y_{t-\ell+1}$. The training set size $n$ should be considerably larger than $\ell$. By Remark 3.2, we can do this by thresholding appropriately the quantity $\sum_{r=0}^{\ell-1}c_{r}Y_{t-r}$, where
@@
4.4 Prediction Results
Table 4.1 summarizes the results from our prediction experiments. We sought to predict exceedances of the $p$th marginal quantile $h$ steps ahead, for $p=0.90,0.95,0.99$ and $h=1$ and $6$ (see also Table A.3 for $h=12$ and $18$). The number of observations used to compute predictions from FARIMA models was 168, the number of hours in one week; this value was also used as the order of the AR models that we fit. The True Skill Statistic (TSS) is commonly used to evaluate solar flare forecasting methods; see Remark A.7 for more information about it. Since the patterns in the TSS section of Table 4.1 are very similar to those in the precision section, we shall only comment on the latter. In each row, in the section for each metric, the largest number is in bold.